# OptMerge: Unifying Multimodal LLM Capabilities and Modalities via Model Merging

### v2

@@ 6 Conclusion
Model merging aims to combine multiple expert models into a single model without requiring data. In this paper, we introduce the model merging benchmark with detailed categorization of MLLM capabilities, and explore how model merging can effectively combine different modalities of MLLMs. We further propose a novel merging method that effectively removes noise from task vectors and improves the robustness of merged vector optimization. Our results demonstrate that model merging potentially surpasses mixture training, serving as a way for omni model alignment, while offering a scalable solution for developing MLLMs with reduced computational cost and time.
@@ Appendix A Theoretical Proofs
Assumption A.1 (Lipschitz continuous).
The loss function used for task $i$ is $\mathcal{C}_{i}$-Lipschitz continuous; that is, for any $\Theta$ and $\Theta^{\prime}$: $\left\|\mathcal{L}_{i}(\Theta)-\mathcal{L}_{i}(\Theta^{\prime})\right\|\leq\mathcal{C}_{i}\left\|\Theta-\Theta^{\prime}\right\|.$ (5)
Lipschitz continuity of the loss function represents a standard assumption in optimization theory and is widely employed to derive generalization bounds and conduct stability analysis. Consequently, we adopt this assumption to analyze the effect of the loss function in our work.
Assumption A.2 (Upper bound of second moment).
Denote the stochastic gradient in step $t$ during the optimization process as $g_{t}$. There exists a constant $\mathcal{G}>0$ such that: $\mathbb{E}\left[\|g_{t}(\Theta)\|^{2}\right]\leq\mathcal{G}^{2}.$ (6)
Assumption A.2 has been formulated in various ways throughout the literature. For convex conditions, Gower et al. (2019) proposed the bound $4\mathcal{C}\bigl(\mathcal{L}(\Theta)-\mathcal{L}(\Theta_{*})\bigr)+2\sigma^{2}$, where $\Theta_{*}$ is the global minimizer and $\sigma^{2}\triangleq\mathbb{E}\left[\left\|g(\Theta_{*})\right\|^{2}\right]$. Similarly, for non-convex settings, Khaled & Richtárik (2023) introduced the expected smoothness assumption as $2A\bigl(\mathcal{L}(\Theta)-\mathcal{L}^{\inf}\bigr)+B\,\|\nabla\mathcal{L}(\Theta)\|^{2}+C$, for some $A,B,C\geq 0$ and $\forall\Theta\in\mathbb{R}^{d}$.
Lemma A.3 (Upper bound of loss affection).
Denote the merged task vector as $\bm{\tau}_{m}=\sum_{j}\alpha_{j}\bm{\tau}_{j}$, and the task vector of task $i$ as $\bm{\tau}_{i}$. Let $\mathcal{L}_{i}$ be the loss for task ii. The difference in loss is bounded by the following: $\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\textstyle\sum_{j}\Psi_{j}^{i}\cdot\left\|\bm{\tau}_{j}\right\|,$ (7) where $\Psi_{j}^{i}$ is the coefficient capturing how sensitive task $i$’s loss is to parameter updates from task $j$ defined as: $\Psi_{j}^{i}=\begin{cases}\mathcal{C}_{i}\cdot|\alpha_{j}|,&\text{if }j\neq i\\ \mathcal{C}_{i}\cdot|1-\alpha_{j}|,&\text{if }j=i.\end{cases}$ (8)
@@
Under assumption that each $\mathcal{L}_{i}$ is $C_{i}$-Lipschitz, we have:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|$ $\leq\mathcal{C}_{i}\cdot\left\|\bm{\tau}_{m}-\bm{\tau}_{i}\right\|$ (9)
$=\mathcal{C}_{i}\cdot\textstyle\left\|\sum_{j}\alpha_{j}\,\bm{\tau}_{j}-\bm{\tau}_{i}\right\|$ (10)
$\leq\mathcal{C}_{i}\cdot(\sum_{j\neq i}\alpha_{j}\cdot\left\|\bm{\tau}_{j}\right\|$
$\quad+\,|1-\alpha_{i}|\cdot\|\bm{\tau}_{i}\|).$ (11)
We define the coefficient $\Psi^{i}_{j}$ for task $i$ as follows:
$\Psi_{j}^{i}=\begin{cases}\mathcal{C}_{i}\cdot|\alpha_{j}|,&\text{if }j\neq i\\ \mathcal{C}_{i}\cdot|1-\alpha_{j}|,&\text{if }j=i.\end{cases}$ (12)
Applying Eq.˜12 to Eq.˜11, we obtain:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\textstyle\sum_{j}\Psi_{j}^{i}\cdot\left\|\bm{\tau}_{j}\right\|.$ (13)
Theorem A.4.
Let $\bm{\theta}_{0}$ be the initial parameter. For each task $i$, after $T$ gradient steps with constant learning rate $\eta$, we denote the task vector as $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$. Considering that most methods can be viewed as extensions of linear combinations of task vectors, let $\bm{\tau}_{m}=\sum_{j}\alpha_{j}\bm{\tau}_{j}$ denote the merged vector. The loss on task $i$ is denoted by $\mathcal{L}_{i}(\Theta)$, which is $\mathcal{C}_{i}$-Lipschitz continuous. Assuming a constraint on the second moment of the gradient, then: $\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\mathcal{O}(\eta T)$ (14) This indicates that both the learning rate and iterations influence model merging results.
@@
Considering the process of gradient update:
$\bm{\tau}_{j}=\textstyle\sum_{t=0}^{T-1}-\eta\cdot g_{t}(\Theta),$ (15)
we have:
$\mathbb{E}[\|\bm{\tau}_{j}\|^{2}]$ $=\mathbb{E}[\|\textstyle\sum_{t=0}^{T-1}-\eta\cdot g_{t}(\Theta)\|^{2}]$ (16)
$=\mathbb{E}[\eta^{2}\|\textstyle\sum_{t=0}^{T-1}g_{t}(\Theta)\|^{2}]$ (17)
$\leq\textstyle\eta^{2}\cdot T\cdot\sum_{t=0}^{T-1}\mathbb{E}\left[\|g_{t}(\Theta)\|^{2}\right]$ (18)
$\leq\eta^{2}\cdot T^{2}\cdot\mathcal{G}^{2}.$ (19)
Since $\mathbb{E}[X^{2}]\geq\mathbb{E}^{2}[X]$, we can derive:
$\mathbb{E}\left[\|\bm{\tau}_{j}\|\right]$ $\leq\sqrt{\mathbb{E}\left[\|\bm{\tau}_{j}\|^{2}\right]}$ (20)
$=\eta\cdot T\cdot\mathcal{G}.$ (21)
Applying Eq.˜20 to Lemma A.3, we obtain:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\underbrace{(\textstyle\sum_{j}\Psi_{j}^{i})\cdot\mathcal{G}}_{\text{coefficient}}\cdot\eta T.$ (22)
Therefore, we can summarize as follows:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\mathcal{O}(\eta T).$ (23)
@@ B.1 Fine-Tuning Influence on Model Merging
To demonstrate the sensitivity of model merging to task vectors $\bm{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base), we conduct experiments using the standard CLIP-ViT merging benchmark (Ilharco et al., 2023), following the fine-tuning setup of FusionBench (Tang et al., 2024a). We train with Adam (learning rate 1e-5) for 4,000 steps with a batch size of 32. Models are saved every 500 iterations and evaluated for accuracy on the test dataset, as illustrated in Fig.˜5. Across eight tasks, convergence typically occurs around 3,000 steps.
@@ Appendix C Implementation Details
We follow prior work (Chen et al., 2024a) by pairing Vicuna-7B-v1.5 with modality-specific encoders and connectors. Separate models are trained using bi-modal data across three modalities: vision, audio, and video. Additional details are presented in Table˜8. The training approach consists of two phases: an alignment stage where only connector parameters are trainable, and a fine-tuning stage where we tune all connector and language model parameters. During fine-tuning, we apply LoRA with a rank of 128 across all linear modules within the LLM. For our merging strategy, we preserve each modality’s unique encoder and connector components while merging only the language model parameters, enabling the model to process inputs from all three modalities simultaneously.
@@
Following (Chen et al., 2024c), we collect a broader range of domain-specific data, divided into VQA, Geometry, Chart, OCR, and Grounding tasks. Specifically, for grounding tasks, we map coordinates to the [0,1000) range and add the special token notation <|box_start|><|box_end|> (Wang et al., 2024b). During Qwen2-VL-Base fine-tuning, we observe that Chinese datasets consistently degraded performance, possibly due to lower data quality or because evaluation benchmarks are primarily in English. Consequently, we use only English datasets for instruction tuning of Qwen2-VL-Base. InternVL2.5-Instruct, already possessing multilingual instruction-following capabilities, is fine-tuned using all available data.
@@
We carefully select specialized datasets to evaluate distinct abilities across tasks. (i) For VQA, we utilize VizWiz (Gurari et al., 2018) and GQA (Hudson & Manning, 2019) to assess general visual question answering proficiency. (ii) For Geometry, we incorporate multiple challenging subsets: “geometry reasoning”, “algebraic reasoning” and “geometry problem solving” from MathVista (Lu et al., 2024a), complemented by “metric geometry - angle”, “metric geometry - area”, “metric geometry - length” and “solid geometry” from MATH-Vision (Wang et al., 2024a). (iii) For Chart, we employ ChartQA (Masry et al., 2022), which tests reasoning and interpretation ability with charts and graphs. (iv) For OCR, our evaluation suite includes TextVQA (Singh et al., 2019) and OCRVQA (Mishra et al., 2019). (v) For Grounding, we implement referring expression comprehension using RefCOCO (Kazemzadeh et al., 2014), RefCOCO+ (Kazemzadeh et al., 2014), and RefCOCOg (Mao et al., 2016), which require models to identify specific objects in images based on natural language descriptions. All evaluation results are obtained using the VLMEvalKit (Duan et al., 2024) and LMMs-Eval (Zhang et al., 2024) libraries under the same settings to ensure fair comparison. When evaluating MathVista and MATH-Vision benchmarks, we utilize the GPT-4o-mini API to extract answers from the output.

### v3

@@ 6 Conclusion
Model merging aims to combine multiple expert models into a single model without requiring data. In this paper, we introduce the model merging benchmark with detailed categorization of MLLM capabilities, and explore how model merging can effectively combine different modalities of MLLMs. We further propose a novel merging method that effectively removes noise from task vectors and improves the robustness of merged vector optimization. Our results demonstrate that model merging potentially surpasses mixture training, serving as a way for omni-model alignment, while offering a scalable solution for developing MLLMs with reduced computational cost and time.
Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant(KJZD20230923115106012,KJZD20230923114916032,GJHZ20240218113604008). This work is supported by National Key R&D Projects (NO. 2024YFC3307100), NSFC Grant (No. 62576364, No. U2541229), Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO. JCYJ20241202124430041). Dr Tao’s research is partially supported by NTU RSR and Start Up Grants.
@@ Appendix A Theoretical Proofs
Notation and setting
Tasks and losses.
For tasks $i$, let the loss be $\mathcal{L}_{i}:\mathbb{R}^{d}\to\mathbb{R}$ evaluated at parameters $\bm{\Theta}\in\mathbb{R}^{d}$.
Task vectors.
For task $i$, after $T$ steps of (deterministic) gradient descent (GD) with fixed step size $\eta>0$ starting from a common initialization $\bm{\Theta}$, the task vector is
$\bm{\tau}_{i}:=-\eta\sum_{t=0}^{T-1}\nabla\mathcal{L}_{i}(\bm{\Theta}^{(i)}_{t})$.
Merged update.
Let the merged vector be
$\bm{\tau}_{m}:=\sum_{j=1}^{m}\alpha_{j}\bm{\tau}_{j},$
with nonnegative weights $\alpha_{j}\geq 0$. We study the loss of task $i$ at the merged point $\bm{\Theta}+\bm{\tau}_{m}$.
Norm and inner product.
$\left\lVert\cdot\right\rVert$ is the Euclidean norm and $\left\langle\cdot,\cdot\right\rangle$ the Euclidean inner product. For nonzero vectors $u,v$, $\cos(u,v):=\left\langle u,v\right\rangle/(\left\lVert u\right\rVert\left\lVert v\right\rVert)$.
Assumption A.1 (L-smoothness).
Each $\mathcal{L}_{i}$ has $L$-Lipschitz continuous gradients; that is, for all $\bm{\Theta},\bm{\Theta}^{\prime}$: $\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta})-\nabla\mathcal{L}_{i}(\bm{\Theta}^{\prime})\right\rVert\leq L\left\lVert\bm{\Theta}-\bm{\Theta}^{\prime}\right\rVert.$ Equivalently, for any $\Delta$: $\mathcal{L}_{i}(\bm{\Theta}+\Delta)\leq\mathcal{L}_{i}(\bm{\Theta})+\left\langle\nabla\mathcal{L}_{i}(\bm{\Theta}),\Delta\right\rangle+\frac{L}{2}\left\lVert\Delta\right\rVert^{2}.$
Assumption A.2 (Polyak–Łojasiewicz (PL) condition).
Each $\mathcal{L}_{i}$ satisfies, for some $\mu>0$: $\frac{1}{2}\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta})\right\rVert^{2}\geq\mu\big(\mathcal{L}_{i}(\bm{\Theta})-\mathcal{L}_{i}^{*}\big),$ where $\mathcal{L}_{i}^{*}:=\inf_{\bm{\Theta}}\mathcal{L}_{i}(\bm{\Theta})$.
Assumption A.3 (Directional similarity).
For each $i$: $\cos\big(-\nabla\mathcal{L}_{i}(\bm{\Theta}),\bm{\tau}_{i}\big)\geq\kappa,\quad\kappa\in(0,1],$ equivalently: $\left\langle\nabla\mathcal{L}_{i}(\bm{\Theta}),\bm{\tau}_{i}\right\rangle\leq-\kappa\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta})\right\rVert\,\left\lVert\bm{\tau}_{i}\right\rVert.$
This ensures that the update is indeed a descent direction for task $i$, with alignment quantified by $\kappa$.
Assumption A.4 (Approximate orthogonality).
For all $i\neq j$: $\cos(\bm{\tau}_{i},\bm{\tau}_{j})\leq\varepsilon,\quad\varepsilon\in[0,1).$
Prior works (Ilharco et al., 2023; Ortiz-Jimenez et al., 2023) show that task vectors are nearly orthogonal, which helps explain the success of model merging. This likely reflects a general property of high-dimensional spaces: independent directions tend to be almost orthogonal. A small $\varepsilon$ means that tasks are nearly orthogonal in update space, reducing negative transfer effects.
Assumption A.5 (Bounded gradients).
There exists $G>0$ such that for all $i$ and all $\bm{\Theta}$ considered: $\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta})\right\rVert\leq G.$
This assumption is widely adopted in the optimization literature (Gower et al., 2019; Khaled and Richtárik, 2023), where similar boundedness conditions are imposed to control the variance of stochastic gradients and to derive finite-step convergence rates.
Lemma A.6 (Cross-task cosine leakage).
Under Assumptions A.3A.4, with $\nabla\mathcal{L}_{i}(\bm{\Theta})\neq\mathbf{0}$ and $\bm{\tau}_{j}\neq\mathbf{0}$ to ensure the cosine is well-defined, for $i\neq j$ we have $\big|\cos\big(\nabla\mathcal{L}_{i}(\bm{\Theta}),\bm{\tau}_{j}\big)\big|\leq\delta,\quad\delta:=\kappa\varepsilon+\sqrt{1-\kappa^{2}}\,\sqrt{1-\varepsilon^{2}}.$
Proof sketch.
Normalize $u=-\nabla\mathcal{L}_{i}/\left\lVert\nabla\mathcal{L}_{i}\right\rVert$, $v_{i}=\bm{\tau}_{i}/\left\lVert\bm{\tau}_{i}\right\rVert$, $v_{j}=\bm{\tau}_{j}/\left\lVert\bm{\tau}_{j}\right\rVert$. Use Assumption A.3 to get $\left\langle u,v_{i}\right\rangle\geq\kappa$ and Assumption A.4 to get $\left\langle v_{i},v_{j}\right\rangle\leq\varepsilon$, then decompose $u$ and $v_{j}$ along $v_{i}$ and its orthogonal complement and apply Cauchy–Schwarz. ∎
Lemma A.7 (PL convergence under GD).
Under Assumptions A.1A.2 and $\eta\in(0,1/L]$, the GD iterates for task $i$ satisfy $\mathcal{L}_{i}(\bm{\Theta}_{T})-\mathcal{L}_{i}^{*}\leq(1-\eta\mu)^{T}\big(\mathcal{L}_{i}(\bm{\Theta}_{0})-\mathcal{L}_{i}^{*}\big).$
@@
Let one step of GD be $\bm{\Theta}_{t+1}=\bm{\Theta}_{t}-\eta\nabla\mathcal{L}_{i}(\bm{\Theta}_{t})$. By L-smoothness, for any $x,y$,
$\mathcal{L}_{i}(y)\leq\mathcal{L}_{i}(x)+\left\langle\nabla\mathcal{L}_{i}(x),y-x\right\rangle+\frac{L}{2}\left\lVert y-x\right\rVert^{2}.$
Plug $x=\bm{\Theta}_{t}$, $y=\bm{\Theta}_{t+1}$:
$\mathcal{L}_{i}(\bm{\Theta}_{t+1})\leq\mathcal{L}_{i}(\bm{\Theta}_{t})-\eta\left(1-\frac{L\eta}{2}\right)\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta}_{t})\right\rVert^{2}.$
Since $\eta\leq 1/L$, we have $1-\frac{L\eta}{2}\geq\frac{1}{2}$, thus
$\mathcal{L}_{i}(\bm{\Theta}_{t+1})\leq\mathcal{L}_{i}(\bm{\Theta}_{t})-\frac{\eta}{2}\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta}_{t})\right\rVert^{2}.$
Apply the PL inequality:
$\frac{1}{2}\left\lVert\nabla\mathcal{L}_{i}(\bm{\Theta}_{t})\right\rVert^{2}\geq\mu\big(\mathcal{L}_{i}(\bm{\Theta}_{t})-\mathcal{L}_{i}^{*}\big),$
to get
$\mathcal{L}_{i}(\bm{\Theta}_{t+1})-\mathcal{L}_{i}^{*}\leq(1-\eta\mu)\big(\mathcal{L}_{i}(\bm{\Theta}_{t})-\mathcal{L}_{i}^{*}\big).$
Unrolling the recursion yields the claim. ∎
Lemma A.8 (Task vector norm).
If $\bm{\tau}_{j}=-\eta\sum_{t=0}^{T-1}\nabla\mathcal{L}_{j}(\bm{\Theta}^{(j)}_{t})$ and $\left\lVert\nabla\mathcal{L}_{j}(\bm{\Theta}^{(j)}_{t})\right\rVert\leq G$ for all $t$, then $\left\lVert\bm{\tau}_{j}\right\rVert\leq\eta TG.$
@@
By the triangle inequality,
$\left\lVert\bm{\tau}_{j}\right\rVert\leq\eta\sum_{t=0}^{T-1}\left\lVert\nabla\mathcal{L}_{j}(\bm{\Theta}^{(j)}_{t})\right\rVert\leq\eta\sum_{t=0}^{T-1}G=\eta TG.$
Lemma A.9 (Inner-product lower bound).
Under Assumptions A.1A.2 and $\eta\in(0,1/L]$, $\left\langle\nabla\mathcal{L}_{i}(\bm{\Theta}),\bm{\tau}_{i}\right\rangle\geq-\Big(1-(1-\eta\mu)^{T}\Big)\big(\mathcal{L}_{i}(\bm{\Theta})-\mathcal{L}_{i}^{*}\big)-\frac{L}{2}\left\lVert\bm{\tau}_{i}\right\rVert^{2}.$
Proof.
Apply the $L$-smooth upper bound with $\Delta=\bm{\tau}_{i}$ and rearrange; then use Lemma A.7 to bound $\mathcal{L}_{i}(\bm{\Theta}+\bm{\tau}_{i})-\mathcal{L}_{i}(\bm{\Theta})$. ∎
Theorem A.10 (Finite-step bound).
Consider task $i$ trained for $T$ iterations of gradient descent with a fixed step size $\eta\in(0,1/L]$. Let $\gamma:=1-\eta\mu\in(0,1)$ denote the PL convergence factor. Then the merged update $\bm{\tau}_{m}:=\sum_{j=1}^{m}\alpha_{j}\bm{\tau}_{j}$ satisfies $\mathcal{L}_{i}(\bm{\Theta}+\bm{\tau}_{m})\leq C_{i}+\mathcal{O}(\gamma^{T})+\mathcal{O}(\delta\,\eta T)+\mathcal{O}(\eta^{2}T^{2}),$ where $\mathcal{O}(\gamma^{T})$ is the residual error from incomplete convergence on task $i$, $\mathcal{O}(\delta\,\eta T)$ is the cross-task interference term, and $\mathcal{O}(\eta^{2}T^{2})$ is the curvature term from $L$-smoothness.
Proof.
Define the $\eta,T$-independent constant
$C_{i}:=\mathcal{L}_{i}(\bm{\Theta})-\alpha_{i}\big(\mathcal{L}_{i}(\bm{\Theta})-\mathcal{L}_{i}^{*}\big).$
By $L$-smoothness,
$\mathcal{L}_{i}(\bm{\Theta}+\bm{\tau}_{m})\leq\mathcal{L}_{i}(\bm{\Theta})+\left\langle\nabla\mathcal{L}_{i}(\bm{\Theta}),\bm{\tau}_{m}\right\rangle+\frac{L}{2}\left\lVert\bm{\tau}_{m}\right\rVert^{2}.$
Decomposing the inner product yields
$\left\langle\nabla\mathcal{L}_{i},\bm{\tau}_{m}\right\rangle=\alpha_{i}\left\langle\nabla\mathcal{L}_{i},\bm{\tau}_{i}\right\rangle+\sum_{j\neq i}\alpha_{j}\left\langle\nabla\mathcal{L}_{i},\bm{\tau}_{j}\right\rangle.$
For the self term, Lemma A.9 implies a constant part absorbed into $C_{i}$ and a residual term of order $\mathcal{O}(\gamma^{T})$, plus a curvature correction $\mathcal{O}(\eta^{2}T^{2})$ via Lemma A.8. For the cross terms, Lemma A.6 and Assumption A.5 give
$\big|\left\langle\nabla\mathcal{L}_{i},\bm{\tau}_{j}\right\rangle\big|\leq\delta\,\eta T\,G^{2},$
so the sum over $j\neq i$ is $\mathcal{O}(\delta\,\eta T)$. Finally, $\left\lVert\bm{\tau}_{m}\right\rVert\leq\eta TG\sum_{j}\alpha_{j}$ implies the smoothness term is $\mathcal{O}(\eta^{2}T^{2})$. Combining all contributions yields the stated bound. ∎
Theorem A.11 (Bound in the near-convergence regime).
Suppose the residual PL error after $T$ steps is below a given tolerance $\zeta>0$: $(1-\eta\mu)^{T}\big(\mathcal{L}_{i}(\bm{\Theta})-\mathcal{L}_{i}^{*}\big)\leq\zeta.$ Equivalently, $T\geq\frac{\ln\!\big((\mathcal{L}_{i}(\bm{\Theta})-\mathcal{L}_{i}^{*})/\zeta\big)}{-\ln(1-\eta\mu)}.$ Then the merged loss satisfies $\mathcal{L}_{i}(\bm{\Theta}+\bm{\tau}_{m})\leq C_{i}+\mathcal{O}(\zeta)+\mathcal{O}(\delta\,\eta T)+\mathcal{O}(\eta^{2}T^{2}),$ with the same $C_{i}$ as in Theorem A.10.
Proof.
Starting from Theorem A.10, replace the residual term $\mathcal{O}(\gamma^{T})$ by $\mathcal{O}(\zeta)$ using the near-convergence assumption. The cross-task and curvature terms remain unchanged. ∎
Remark: When the learning rate is fixed and the model has not yet converged, the improvement from finetuning on the target task (captured by $1-\gamma^{T}$) typically outweighs the influence of other task vectors, especially when these vectors are close to orthogonal (small $\varepsilon$, hence small $\delta$). In this stage, merging different task updates remains stable and can be beneficial.
However, as training approaches convergence, the potential negative impact from other task vectors becomes more significant. Even if each single-task loss continues to decrease, the merged model’s loss can worsen due to accumulated cross-task interference, which grows linearly in $T$ as $\mathcal{O}(\delta\,\eta T)$, and curvature effects from $L$-smoothness, which grow quadratically as $\mathcal{O}(\eta^{2}T^{2})$. This means that over-training on individual tasks can harm the quality of the merged model.
In the convergence regime, a fixed learning rate can lead to increased norms of task vectors. While the single-task losses may remain similar, the larger norms amplify interference and curvature errors, further degrading merged performance. Once the PL error $(1-\eta\mu)^{T}(\mathcal{L}_{i}-\mathcal{L}_{i}^{*})$ falls below a tolerance $\zeta$, the main residual terms are the interference and curvature contributions. Reducing directional leakage (small $\delta$) and controlling the product $\eta T$ are therefore essential for high-quality merging.
@@ B.1 Fine-Tuning Influence on Model Merging
To demonstrate the sensitivity of model merging to task vectors $\bm{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base), we conduct experiments using the standard CLIP-ViT merging benchmark, following the fine-tuning setup of FusionBench (Tang et al., 2024a). We train with Adam (learning rate 1e-5) for 4,000 steps with a batch size of 32. Models are saved every 500 iterations and evaluated for accuracy on the test dataset, as illustrated in Fig.˜5. Across eight tasks, convergence typically occurs around 3,000 steps.
@@ Appendix C Implementation Details
We follow prior work (Chen et al., 2024a) by pairing Vicuna-7B-v1.5 with modality-specific encoders and connectors. Separate models are trained using bi-modal data across three modalities: vision, audio, and video. Additional details are presented in Table˜11. The training approach consists of two phases: an alignment stage where only connector parameters are trainable, and a fine-tuning stage where we tune all connector and language model parameters. During fine-tuning, we apply LoRA with a rank of 128 across all linear modules within the LLM. For our merging strategy, we preserve each modality’s unique encoder and connector components while merging only the language model parameters, enabling the model to process inputs from all three modalities simultaneously.
@@
Following (Chen et al., 2024c), we collect a broader range of domain-specific data, divided into VQA, Geometry, Chart, OCR, and Grounding tasks. For each dataset, we use only the training split, containing question, answer, and image. Samples exceeding 8192 tokens in combined question–answer length or with corrupted images are removed. The remaining data are converted into the ShareGPT instruction-tuning format. Tasks (e.g., VQA, OCR) are trained separately, so cross-task balancing is unnecessary. We collect at least 100k public samples per task to ensure diversity, following common practice for fine-tuning models in the 1B–7B parameter range. Specifically, for grounding tasks, we map coordinates to the [0,1000) range and add the special token notation <|box_start|><|box_end|> (Wang et al., 2024b). During Qwen2-VL-Base fine-tuning, we observe that Chinese datasets consistently degraded performance, possibly due to lower data quality or because evaluation benchmarks are primarily in English. Consequently, we use only English datasets for instruction tuning of Qwen2-VL-Base. InternVL2.5-Instruct, already possessing multilingual instruction-following capabilities, is fine-tuned using all available data.
@@
We carefully select specialized datasets to evaluate distinct abilities across tasks. (i) For VQA, we utilize VizWiz (Gurari et al., 2018) and GQA (Hudson and Manning, 2019) to assess general visual question answering proficiency. (ii) For Geometry, we incorporate multiple challenging subsets: “geometry reasoning”, “algebraic reasoning” and “geometry problem solving” from MathVista (Lu et al., 2024a), complemented by “metric geometry - angle”, “metric geometry - area”, “metric geometry - length” and “solid geometry” from MATH-Vision (Wang et al., 2024a). (iii) For Chart, we employ ChartQA (Masry et al., 2022), which tests reasoning and interpretation ability with charts and graphs. (iv) For OCR, our evaluation suite includes TextVQA (Singh et al., 2019) and OCRVQA (Mishra et al., 2019). (v) For Grounding, we implement referring expression comprehension using RefCOCO (Kazemzadeh et al., 2014), RefCOCO+ (Kazemzadeh et al., 2014), and RefCOCOg (Mao et al., 2016), which require models to identify specific objects in images based on natural language descriptions. All evaluation results are obtained using the VLMEvalKit (Duan et al., 2024) and LMMs-Eval (Zhang et al., 2024) libraries under the same settings to ensure fair comparison. When evaluating MathVista and MATH-Vision benchmarks, we utilize the GPT-4o-mini API to extract answers from the output. The following prompt is used, where {question} denotes the question text and {prediction} denotes the original output from the evaluated model.
Please read the following examples. Then extract the answer from the model response and type it at the end of the prompt.
Hint: Please answer the question requiring an integer answer and provide the final value,
e.g., 1, 2, 3, at the end.
Question: Which number is missing?
Model response: The number missing in the sequence is 14.
Extracted answer: 14
Hint: Please answer the question requiring a floating-point number with one decimal place and provide the final value,
e.g., 1.2, 1.3, 1.4, at the end.
Question: What is the fraction of females facing the camera?
Model response: The fraction of females facing the camera is 0.6,
which means that six out of ten females in the group are facing the camera.
Extracted answer: 0.6
Hint: Please answer the question requiring a floating-point number with two decimal places and provide the final value,
e.g., 1.23, 1.34, 1.45, at the end.
Question: How much money does Luca need to buy a sour apple candy and a butter-scotch candy? (Unit: \$)
Model response: Luca needs \$1.45 to buy a sour apple candy and a butterscotch candy.
Extracted answer: 1.45
Hint: Please answer the question requiring a Python list as an answer and provide the final list,
e.g., [1, 2, 3], [1.2, 1.3, 1.4], at the end.
Question: Between which two years does the line graph saw its maximum peak?
Model response: The line graph saw its maximum peak between 2007 and 2008.
Extracted answer: [2007, 2008]
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
Question: What fraction of the shape is blue?
Choices: (A) 3/11 (B) 8/11 (C) 6/11 (D) 3/5
Model response: The correct answer is (B) 8/11.
Extracted answer: B
{question}
Model response: {prediction}
Extracted answer: